OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1 A. -69x + 10y = -51 B. 69x + 10y = -51 C. 69x + 10y = 51 D. -69x + 10y = 51

jimthompson5910 (jim_thompson5910):

First, you need every number to become a whole number so because we have decimal numbers, we need to change that

jimthompson5910 (jim_thompson5910):

6.9 and 5.1 both have 1 decimal place so multiplying them by 10^1 = 10 will move that decimal 1 spot to the right

jimthompson5910 (jim_thompson5910):

multiply both sides by 10 y = -6.9x + 5.1 10y = 10(-6.9x + 5.1) 10y = -69x + 51 Notice how the decimal numbers are now whole numbers

jimthompson5910 (jim_thompson5910):

The last step is to add 69x to both sides to cancel out the -69x on the right side we do this because we want to move over that 69x